OpenStudy (anonymous):

A combined circuit has two resistors in parallel (10.0 ohms and 14.0 ohms) and another in series (5.0 ohms). If the power source is 9.0 volts, what is the current?

OpenStudy (surry99):

The first step is to draw the circuit. The second is determine the overall equivalent resistance of the circuit.

OpenStudy (anonymous):

Your circuit would be like this:  Here the total current (i) is the current used by resistance R1 and R2, let total resistance is R (which is addition of R1 and R2) and voltage is V \[i=\frac{ V }{ R }\] let's calculate R first, where R1 is the parallel of two resistances(10 and 14 ohm) giving a resultant\[\frac{ 1 }{ \frac{ 1 }{ 10 }+\frac{ 1 }{ 14 } }\] =5.8333 ohm and R2=5 ohm hence total resistance R= R1 + R2 .....(as R1 and R2 are in series) R=5+5.8333 =10.8333 ohm now \[i=\frac{ 9 }{ 10.8333 }\] that is i=0.83077 A or total current i=830.77 mA